OpenStudy (anonymous):

The length of a rectangular porch is (x+7)ft. The area of the porch is (x^2+9x+14)ft^2. Find the width of the porch.

OpenStudy (anonymous):

\[\frac{x^2+9x+14}{x+7}=x+2\]

OpenStudy (anonymous):

is that the answer

Directrix (directrix):

The area of a rectangle is length times width. x^2 + 9x + 14 = (x+7) times width. width = (x^2 + 9x + 14) /(x + 7) I suspect that (x+7) is a factor of (x^2 + 9x + 14) and will see. (x^2 + 9x + 14) = (x + 7)(x + 2) [ (x + 7)(x + 2)] / (x + 7) = (x + 2) The width of the porch is x + 2 .

Directrix (directrix):

* (x + 2) feet